OpenStudy (anonymous):

The height of a cylinder with a fixed radius of 4 cm is increasing at the rate of 2 cm/min. Find the rate of change of the volume of the cylinder (with respect to time) when the height is 14cm.

OpenStudy (anonymous):

what is the formula for the volume of a cylinder ?

OpenStudy (anonymous):

oh i think it is \\[V=\pi r^2h\] which in your case is \[V=16\pi h\] is that right?

OpenStudy (anonymous):

if so that means \[V'=16\pi h'\] and you are told \(h'=4\) plug it in, see what you get

OpenStudy (michele_laino):

from your data, If I call h the height of the cylinder, and \[\alpha \]=2 cm/min the rate of increasing for h, we can write: \[h(t)=h _{0}+\alpha t\] where \[h _{0}\] is the initial height of cylinder, so if I call V the volume of cylinder, we can write: \[V(t)=16\pi h(t)=16 \pi (h _{0}+a t)\] from which, by derivation, I get: \[\frac{ dV }{ dt }=16 \pi \ \alpha \] I think it is the rate of increasing of the volume of the cylinder, which is independent of t